two sides of a pentagon have length (3x-5) cm and three sides of length (2x-1) cm. if the perimeter is 89 cm, what is the length of longest side?

Answer by checkley77(12844)

2(3X-5)+3(2X-1)=89
6X-10+6X-3=89
12X=89+10+3
12X=102
X=102/12
X=8.5 ANS.
3*8.5-5=25.5-5=20.5 ANS. FOR THE LONGEST SIDES.
2*8.5-1=17-1=16 ANS. FOR THE SHORT SIDES.
PROOF:
2*20.5+3*16=89
41+48=89
89=89